# Moving Object Segmentation: All You Need Is SAM (and Flow)

### v1

@@ Abstract
We investigate two models for combining SAM with optical flow that harness the segmentation power of SAM with the ability of flow to discover and group moving objects. In the first model, we adapt SAM to take optical flow, rather than RGB, as an input. In the second, SAM takes RGB as an input, and flow is used as a segmentation prompt. These surprisingly simple methods, without any further modifications, outperform all previous approaches by a considerable margin in both single and multi-object benchmarks. We also extend these frame-level segmentations to sequence-level segmentations that maintain object identity. Again, this simple model outperforms previous methods on multiple video object segmentation benchmarks.
@@
Motion Segmentation Video Object Segmentation Motion-based Object Discovery
@@ 1 Introduction
To this end, we explore two distinct yet simple variants to effectively tailor SAM for motion segmentation. First, we introduce FlowI-SAM (Fig. 1a), an adaption of the original SAM that directly processes optical flow as a three-channel input image for segmentation, where points on a uniform grid are used as prompts. This approach leverages the ability of SAM to accurately segment moving objects against the static background, by exploiting the distinct textures and clear boundaries present in optical flow fields. However, it has less success in scenes where the optical flow arises from multiple interacting objects as the flow only contains limited information for separating them. Second, building on the strong ability of SAM on RGB image segmentation, we propose FlowP-SAM (Fig. 1b) where the input is an RGB frame, and flow guides SAM for moving object segmentation as prompts, produced by a trainable prompt generator. This method effectively leverages the ability of SAM on RGB image segmentation, with flow information acting as a selector of moving objects/regions within a frame. Additionally, we extend these methods from frame-level to sequence-level video segmentation (Fig. 1c) so that object identities are consistent throughout the sequence. We do this by introducing a matching module that auto-regressively chooses whether to select a new object or propagate the old one based on temporal consistency.
@@
    New state-of-the-art performance by a large margin on moving object segmentation benchmarks, including DAVIS, YTVOS, and MoCA, at both frame and sequence levels.
@@ 2 Related Work
Video object segmentation (VOS) is an extensively studied task in computer vision. The objective is to segment the primary object(s) in a video sequence. Numerous benchmarks are developed for evaluating VOS performance, catering to both single-object [31, 18, 29, 16] and multi-object [33, 46] scenarios. Two major protocols are widely explored in VOS research, namely unsupervised [49, 47, 21, 38, 8, 19, 22, 20] and semi-supervised VOS [39, 14, 13, 27, 40, 11, 2, 30, 50, 6]. Notably, the term “unsupervised” exclusively indicates that no groundtruth annotation is used during inference time (i.e., no inference-time supervision). In contrast, the semi-supervised VOS employs first-frame groundtruth annotations to initiate the object tracking and mask propagation in subsequent frames. This paper focuses on unsupervised VOS tasks and utilises motion as a crucial cue for object discovery.
@@
Segment Anything Model (SAM) [12] has demonstrated impressive ability on image segmentation across various scenarios. It was trained on the SA-1B datasets with over one billion self-labelled masks and $11$ million images. Such large-scale training renders it a strong zero-shot generalisability to unseen domains. Many works adapt the SAM model to perform different tasks, such as tracking [7], change detection [56], and 3D segmentation [3]. Some other works extend SAM towards more promptability [58], more efficient models [51, 55, 45], or more domains [23, 42, 34, 4]. Our work extends the robust SAM framework to moving object segmentation by adapting it with optical flow inputs and flow prompts.
@@ 4 Frame-Level Segmentation I: Flow as Input
To adapt SAM for this new task, we formulate FlowI-SAM ($\Phi_{\text{{\tt FlowI-SAM}}}$) by finetuning it on optical Flow Inputs, and re-purpose the original IoU prediction head to instead predict the fIoU, as illustrated in Fig. 2b.
Flow Inputs with Multiple Frame Gaps. To mitigate the effect of noisy optical flow, i.e. complicated flow fields due to stationary parts, articulated motion, and object interactions, etc., we consider multiple flow inputs $\{F_{t,g}\}$ with different frame gaps (e.g., $g\in\{(1,$-1),(2,1),(2,1 ) , ( 2 ,-2)}2)\}2 ) }) for both training and evaluation stages. These multi-gap flow inputs are independently processed by the image encoder to obtain dense spatial features $\{d_{t,g}\}$ at a lower resolution $h\times w$, which are then combined by averaging the spatial feature maps across different flow gaps, i.e., $d_{t}=\text{Average}_{g}(\{d_{t,g}\})\in\mathbb{R}^{h\times w\times d}$.
@@ 5 Frame-Level Segmentation II: Flow as Prompt
In this section, we adapt SAM to video object segmentation by processing RGB frames, with optical flow as a prompt. We term this frame-level segmentation architecture FlowP-SAM for Flow as Prompt SAM. As shown in Fig. 3b, FlowP-SAM encompasses two major modules, namely the flow prompt generator and the segmentation module. The flow prompt generator takes optical flow as inputs, and produces flow prompts that can be used as supplemental queries to infer frame-level segmentation masks ${M^{i}_{t}}$ from RGB inputs $I_{t}$. Formally,
@@
where $s_{\text{MOS},t}^{i}$ indicates the moving object score (MOS) predicted by the flow prompt generator, while $s_{\text{fIoU},t}^{i}$ denotes the foreground object IoU (fIoU) estimated by the segmentation module. Specifically, MOS specifies whether the input point prompt belongs to moving objects ($s_{\text{MOS},t}^{i}=1$) or the stationary region ($s_{\text{MOS},t}^{i}=0$). On the other hand, fIoU follows the same formulation as in the FlowI-SAM, i.e., predicting IoUs for foreground objects and yielding 00 for background regions.
@@
Flow Prompt Generator consists of (i) a frozen SAM image encoder, where the dense spatial features are extracted from optical flow inputs at different frame gaps, followed by an averaging across frame gaps; (ii) a flow transformer, with the detailed architecture depicted in Fig. 3c, where we first concatenate the input point prompt (i.e., a positional embedding) with learnable flow prompts and moving object score (MOS) tokens to form queries. These queries then iteratively attend to dense flow features in a lightweight transformer decoder. There are two outputs from the flow prompt generator, namely, the refined flow prompts, and an MOS token, which is subsequently processed by an MLP-based head to yield a final moving object score.
Segmentation Module. The overall structure of the segmentation module resembles the original SAM, except for two adaptions: (i) The IoU-prediction head is re-purposed to predict foreground object scores (fIoU) (same as the FlowI-SAM); (ii) The outputs tokens from flow prompt generator are injected as additional query inputs.
@@
FlowP-SAM Training. We train FlowP-SAM in an end-to-end fashion while keeping the SAM pre-trained prompt encoder and image encoders frozen. The flow transformer is trained from scratch, and the mask decoder in the segmentation module is finetuned on top of the pre-trained SAM. The overall loss is:
@@ 6 Sequence-level Mask Association
We then consider three sets of masks: (i) the warped masks $\{\mathscr{M}^{i}_{t\leftarrow t-1}\}$; (ii) the frame-level predictions $\{{M}^{i}_{t}\}$ at frame $t$; (iii) the frame-level predictions from neighboring frames (with $\Delta{t}$ gap) after aligning them to the current frame using optical flow (i.e., $\{{M}^{i}_{t\leftarrow t+\Delta{t}}\}$). For each pair of mask sets, we perform a pairwise Hungarian matching based on the IoU score, resulting in three pairings in total. The Hungarian matched pairs can then reflect the temporal consistency across these predictions based on the transitivity principle, i.e. if object $i$ in (i) matches with object $j$ in (ii) and object $k$ in (iii), then the latter two objects must also match with each other. If such transitivity holds, we set the consistency score $c_{i}=1$, and $c_{i}=0$ otherwise.
@@
Note that we do this separately for each object $i\in N$, so each object gets updated or propagated independently. We lastly layer all the objects back together (to its original order) and remove any overlaps to obtain the final sequence-level predictions.
@@ 7.1 Datasets
Summary of Evaluation Datasets. To investigate the role of motion in object discovery and segmentation, we adopt all aforementioned benchmarks, which consist of only predominantly moving object sequences. Notably, for the evaluation of FlowI-SAM, we exclude the class-labelled DAVIS2017 dataset, as commonly moving objects cannot be separated solely based on motion cues.
Training Datasets. To adapt the RGB pre-trained SAM for moving object discovery and motion segmentation, we train both FlowI-SAM and FlowP-SAM first on the synthetic dataset introduced by [43], as described below, and then on real-world video datasets, including DAVIS2016, DAVIS2017, and DAVIS2017-motion.
@@ 7.3 Implementation Details
Flow Computation. We adopt an off-the-shelf method (RAFT [37]) to estimate optical flow with multiple frame gaps at (1,(1,( 1 ,-1)1)1 ) and (2,(2,( 2 ,-2)2)2 ), except for YTVOS18-m and FBMS-59, where higher frame gaps at (3,(3,( 3 ,-3)3)3 ) and (6,(6,( 6 ,-6)6)6 ) are used to compensate for slow motion.
@@ 7.4 Ablation Study
We present a series of ablation studies on key designs in the per-frame FlowI-SAM
and FlowP-SAM models, as well as on our sequence-level method. For a more detailed ablation analysis, we refer the reader to Appendix 0.B.
$1,$$1$ $64.5$ $78.0$
$2,$$2$ $65.3$ $77.7$
$1,$-$1,2,$-$2$ $65.7$ $79.1$
Table 1: The frame gaps of optical flow inputs to FlowI-SAM. The SAM ViT-H image encoder is adopted. The results are shown for frame-level predictions.
Taking max $65.0$ $78.2$
Averaging $65.7$ $79.1$
Ablation Studies for FlowI-SAM. Below are the investigations focusing on input optical flows and how the dense flow features are combined.
  • $\bullet$
  • $\bullet$
@@ 7.5 Quantitative Results
In terms of the flow-only segmentation, our FlowI-SAM (with both SAM image encoders) outperforms the previous methods by a large margin (>$10\%$). Regarding the RGB-based segmentation, our FlowP-SAM also achieves state-of-the-art performance, particularly excelling at multi-object benchmarks. By combining these two frame-level predictions (FlowI-SAM+FlowP-SAM), we observe further performance boosts. This suggests the complementary roles of the flow and RGB modalities in frame-level segmentation, particularly when there are multiple moving objects involved. In particular, we show that using both models in tandem by layering FlowI-SAM’s predictions behind that of FlowP-SAM allows the model to fill in on missed predictions (such as motion blur, poor lighting, or small objects).
@@
FlowP-SAM+FlowI-SAM (seq) $\checkmark$ $\checkmark$ $\mathbf{74.7}$ $\mathbf{74.3}$ $\mathbf{71.0}$ $\mathbf{87.7}$ $80.1$ $82.8$ $\mathbf{0.647}$
@@
For RGB-based segmentation, we obtain our sequence-level predictions by  FlowI-SAM+FlowP-SAM. As can be seen from the lower part of Table 5, our method achieves outstanding performance across both single- and multi-object benchmarks.
@@ 7.6 Qualitative Visualisations
In this section, example visualisations are provided across multiple datasets, where we mainly present sequence-level results, owing to their marginal differences from frame-level predictions. More visualisations across different evaluation datasets can be found in Appendix 0.E.
@@
8 Conclusion
In this paper, we focus on moving object segmentation in real-world videos, by incorporating per-frame SAM with motion information (optical flow) in two ways: (i) for flow-only segmentation, we introduce FlowI-SAM that directly takes in optical flow as inputs; (ii) for RGB-based segmentation, we utilise motion information to generate flow prompts as guidance. Both approaches deliver state-of-the-art performance in frame-level segmentation across single- and multi-object benchmarks. Additionally, we develop a frame-wise association method that amalgamates predictions from FlowI-SAM  and  FlowP-SAM , achieving sequence-level segmentation predictions that outperform all prior methods on the DAVIS, YTVOS, and MoCA benchmarks.

### v2

@@ Abstract
We investigate two models for combining SAM with optical flow that harness the segmentation power of SAM with the ability of flow to discover and group moving objects. In the first model, we adapt SAM to take optical flow, rather than RGB, as an input. In the second, SAM takes RGB as an input, and flow is used as a segmentation prompt. These surprisingly simple methods, without any further modifications, outperform all previous approaches by a considerable margin in both single and multi-object benchmarks. We also extend these frame-level segmentations to sequence-level segmentations that maintain object identity. Again, this simple model achieves outstanding performance across multiple moving object segmentation benchmarks.
@@
Moving Object Discovery Video Object Segmentation
@@ 1 Introduction
To this end, we explore two simple variants to effectively tailor SAM for motion segmentation. First, we introduce FlowI-SAM (Fig. 1a), an adaption of the original SAM that directly processes optical flow as a three-channel input image for segmentation, where points on a uniform grid are used as prompts. This approach leverages the ability of SAM to accurately segment moving objects against the static background, by exploiting the distinct textures and clear boundaries present in optical flow fields. However, it has less success in scenes where the optical flow arises from multiple interacting objects as the flow only contains limited information for separating them. Second, building on the strong ability of SAM on RGB image segmentation, we propose FlowP-SAM (Fig. 1b) where the input is an RGB frame, and flow guides SAM for moving object segmentation as prompts, produced by a trainable prompt generator. This method effectively leverages the ability of SAM on RGB image segmentation, with flow information acting as a selector of moving objects/regions within a frame. Additionally, we extend these methods from frame-level to sequence-level video segmentation (Fig. 1c) so that object identities are consistent throughout the sequence. We do this by introducing a matching module that auto-regressively chooses whether to select a new object or propagate the old one based on temporal consistency.
@@
    New state-of-the-art unsupervised video object segmentation performance by a large margin on moving object segmentation benchmarks, including DAVIS16, DAVIS17-m, YTVOS18-m, and MoCA.
@@ 2 Related Work
Video Object Segmentation (VOS) is an extensively studied task in computer vision. The objective is to segment the primary object(s) in a video sequence. Numerous benchmarks are developed for evaluating VOS performance, catering to both single-object [33, 19, 30, 17] and multi-object [35, 50] scenarios. Two major protocols are widely explored in VOS research, namely unsupervised [54, 51, 22, 42, 9, 20, 23, 21] and semi-supervised VOS [43, 15, 14, 28, 44, 12, 3, 31, 55, 7, 32]. Notably, the term “unsupervised” exclusively indicates that no groundtruth annotation is used during inference time (i.e., no inference-time supervision). In contrast, the semi-supervised VOS employs first-frame groundtruth annotations to initiate the object tracking and mask propagation in subsequent frames. This paper focuses on unsupervised VOS and utilises motion as a crucial cue for object discovery.
@@
Segment Anything Model (SAM) [13] has demonstrated impressive ability on image segmentation across various scenarios. It was trained on the SA-1B datasets with over one billion self-labelled masks and $11$ million images. Such large-scale training renders it a strong zero-shot generalisability to unseen domains. Many works adapt the SAM model to perform different tasks, such as tracking [8], change detection [62], and 3D segmentation [4]. Some other works extend SAM towards more efficient models [56, 61, 49], and more domains [24, 46, 37, 5]. However, most studies follow the default prompt options in SAM (i.e., points, boxes, and masks). Recent works [64, 39] have shown that more versatile prompts, including scribbles and visual references, can lead to improvements. In this paper, we explore a novel route that prompts SAM with optical flow and demonstrate its effectiveness for moving object segmentation.
@@ 4 Frame-Level Segmentation I: Flow as Input
To adapt SAM for this new task, we formulate FlowI-SAM ($\Phi_{\text{{\tt FlowI-SAM}}}$) by finetuning it on optical Flow Inputs, and re-purpose the original IoU prediction head to instead predict the fIoU, as illustrated in Fig. 2b. By definition, fIoU is a scalar that measures the “objectness”: fIoU is 00 if the mask belongs to the background, and equal to the IoU between predicted and GT object masks for foreground moving object masks. A high fIoU indicates the predicted mask corresponds to the entire object, while a low fIoU might suggest the mask is erroneous or only captures a small part of the object.
Flow Inputs with Multiple Frame Gaps. To mitigate the effect of noisy optical flow, i.e. complicated flow fields due to stationary parts, articulated motion, and object interactions, etc., we consider multiple flow inputs $\{F_{t,g}\}$ with different frame gaps (e.g., $g\in\{(1,$-1),(2,1),(2,1 ) , ( 2 ,-2)}2)\}2 ) }) for both training and evaluation stages. These multi-gap flow inputs are independently processed by the image encoder to obtain dense spatial features $\{d_{t,g}\}$ at a lower resolution $h\times w$, which are then combined by averaging the spatial feature maps across different flow gaps, i.e., $d_{t}=\text{Average}_{g}(\{d_{t,g}\})\in\mathbb{R}^{h\times w\times d}$. These averaged spatial features are then treated as keys and values in the mask decoder.
@@ 5 Frame-Level Segmentation II: Flow as Prompt
In this section, we adapt SAM for video object segmentation by processing RGB frames, with optical flow as a prompt. We term this frame-level segmentation architecture FlowP-SAM for Flow as Prompt SAM. As shown in Fig. 3b, FlowP-SAM encompasses two major modules, namely the flow prompt generator and the segmentation module. The flow prompt generator takes optical flow as inputs, and produces flow prompts that can be used as supplemental queries to infer frame-level segmentation masks ${M^{i}_{t}}$ from RGB inputs $I_{t}$. Formally,
@@
where $s_{\text{MOS},t}^{i}$ indicates the moving object score (MOS) predicted by the flow prompt generator, while $s_{\text{fIoU},t}^{i}$ denotes the foreground object IoU (fIoU) estimated by the segmentation module. Specifically, MOS measures whether each input point prompt (therefore the resultant mask) is within a moving object region based on observing flow fields. Groundtruth MOS scores are binary (i.e., $\hat{s}_{\text{MOS},t}^{i}=1$ if the point prompt is within GT annotation, and $\hat{s}_{\text{MOS},t}^{i}=0$ otherwise). On the other hand, fIoU follows the same formulation as in FlowI-SAM, i.e., predicting IoUs for foreground objects and yielding 00 for background regions.
@@
Flow Prompt Generator consists of (i) a frozen SAM image encoder, where the dense spatial features are extracted from optical flow inputs at different frame gaps, followed by an averaging across frame gaps; (ii) a flow transformer, with the detailed architecture depicted in Fig. 3c, where we first stack the input point prompt (i.e., a positional embedding) with learnable flow prompts and moving object score (MOS) tokens to form queries. These queries then iteratively attend to dense flow features in a lightweight transformer decoder. There are two outputs from the flow prompt generator, namely, the refined flow prompts, and an MOS token, which is subsequently processed by an MLP-based head to yield a final moving object score.
Segmentation Module has a structure that resembles the original SAM, except for two adaptions: (i) The IoU-prediction head is re-purposed to predict foreground object scores (fIoU) (same as the FlowI-SAM); (ii) The outputs tokens from flow prompt generator are injected as additional query inputs.
@@
FlowP-SAM Training. We train FlowP-SAM in an end-to-end fashion while keeping the SAM pre-trained prompt encoder and image encoders frozen. The flow transformer is trained from scratch,
@@ 6 Sequence-level Mask Association
We then consider three sets of masks: (i) the warped masks $\{\mathscr{M}^{i}_{t\leftarrow t-1}\}$; (ii) the frame-level predictions $\{{M}^{i}_{t}\}$ at frame $t$; and (iii) the frame-level predictions from neighboring frames (with $\Delta{t}$ gap) after aligning them to the current frame using optical flow (i.e., $\{{M}^{i}_{t\leftarrow t+\Delta{t}}\}$). For each pair of mask sets, we perform a pairwise Hungarian matching based on the IoU score, resulting in three pairings in total. The Hungarian matched pairs can then reflect the temporal consistency across these predictions based on the transitivity principle, i.e. if object $i$ in (i) matches with object $j$ in (ii) and object $k$ in (iii), then the latter two objects must also match with each other. If such transitivity holds, we set the consistency score $c_{i}=1$, and $c_{i}=0$ otherwise.
@@
Note, we do this separately for each object $i\in N$, which gets updated or propagated independently. We lastly layer all objects back together (to its original order) and remove any overlaps to obtain the final sequence-level predictions.
@@ 7.1 Datasets
Summary of Evaluation Datasets. To investigate the role of motion in object discovery and segmentation, we adopt all aforementioned benchmarks, which consist of only predominantly moving object sequences. Notably, for the evaluation of FlowI-SAM, we exclude the class-labelled DAVIS17 dataset, as commonly moving objects cannot be separated solely based on motion cues.
Training Datasets. To adapt the RGB pre-trained SAM for moving object discovery and motion segmentation, we train both FlowI-SAM and FlowP-SAM first on the synthetic dataset introduced by [47], as described below, and then on real-world video datasets, including DAVIS16, DAVIS17, and DAVIS17-m.
@@ 7.3 Implementation Details
Flow Computation. We adopt an off-the-shelf method (RAFT [41]) to estimate optical flow with multiple frame gaps at (1,(1,( 1 ,-1)1)1 ) and (2,(2,( 2 ,-2)2)2 ), except for YTVOS18-m and FBMS-59, where higher frame gaps at (3,(3,( 3 ,-3)3)3 ) and (6,(6,( 6 ,-6)6)6 ) are used to compensate for slow motion. Following common practice [51, 47], we convert optical flow into 3-channel RGB format using a standard color wheel [1].
@@ 7.4 Ablation Study
In this section, we present a series of ablation studies on key designs in the per-frame FlowP-SAM model. We refer the reader to Appendix 0.B for a more detailed ablation analysis on the designs in FlowI-SAM and FlowP-SAM models, as well as in our sequence-level method.
@@
Discussion on the effectiveness of MOS and fIoU. As outlined in Sect. 3, the original SAM framework over-segments images into objects, parts, and subparts, which cannot be further distinguished using the default SAM IoU estimations. To adapt this setup for object-only discovery, we propose two new scores, i.e., MOS and fIoU, as new criteria to filter out non-object masks. These scores effectively assess the “objectness” of masks: MOS determines if the predicted masks represent moving objects, while fIoU evaluates whether the masks depict complete objects and are not background segments. The effectiveness of this adaptation is validated in Table 1, where replacing IoU estimations with MOS + fIoU scores leads to noticeable performance boosts.
@@ 7.5 Quantitative Results
For flow-only segmentation, our FlowI-SAM (with both SAM image encoders) outperforms the previous methods by a large margin (>$10\%$). For RGB-based segmentation, our FlowP-SAM also achieves state-of-the-art performance, particularly excelling at multi-object benchmarks. By combining these two frame-level predictions (FlowI-SAM+FlowP-SAM), we observe further performance boosts. This suggests the complementary roles of the flow and RGB modalities in frame-level segmentation, particularly when there are multiple moving objects involved. In particular, we show that using both models in tandem by layering FlowI-SAM’s predictions behind that of FlowP-SAM allows the model to fill in on missed predictions (such as motion blur, poor lighting, or small objects).
@@
UVOSAM [59] $\checkmark$ -- -- $\mathbf{77.5}$- -- --
FlowP-SAM+FlowI-SAM (seq) $\checkmark$ $\checkmark$ $\mathbf{74.7}$ $\mathbf{74.3}$ $71.0$ $\mathbf{87.7}$ $80.1$ $82.8$ $\mathbf{0.647}$
@@
For RGB-based segmentation, we obtain our sequence-level predictions by  FlowI-SAM+FlowP-SAM. As shown in the lower part of Table 3, our method achieves outstanding performance across single- and multi-object benchmarks. Note, DAVIS17 annotations are class-based, which could be unclear and inconsistent with the class-agnostic unsupervised VOS methods. More discussion and results are provided in Appendix 0.D.
@@ 7.6 Qualitative Visualisations
In this section, example visualisations are provided across multiple datasets. We refer more visualisations to Appendix 0.E.
@@
8 Discussion
In this paper, we focus on moving object segmentation in real-world videos, by incorporating per-frame SAM with motion information (optical flow) in two ways: (i) for flow-only segmentation, we introduce FlowI-SAM that directly takes in optical flow as inputs; (ii) for RGB-based segmentation (FlowP-SAM), we utilise motion information to generate flow prompts as guidance. The former (FlowI-SAM) is particularly effective in scenarios with predominant motion and/or where RGB information might introduce confusion, such as in moving object detection and camouflaged object discovery. Additionally, owing to simple texture and a low cross-domain gap in optical flow, it generalises to diverse domains beyond everyday videos. On the other hand, FlowP-SAM focuses on common videos where both RGB and motion are informative and can be utilised to resolve ambiguities for independent objects in common motion.
Both approaches deliver state-of-the-art performance in frame-level segmentation across single- and multi-object benchmarks. Additionally, we develop a frame-wise association method that amalgamates predictions from FlowI-SAM  and  FlowP-SAM, achieving sequence-level segmentation predictions that outperform existing methods on DAVIS16, DAVIS17-m, YTVOS18-m, and MoCA.